Word Problem--
Adult tickets for a play cost $4 and child tickets costs $1. If there were 23 people at a performance and the theater collected $80 from ticket sales, how many children attended the play?

Answer by checkley71(8403)

4A+1C=80
A+C=23 OR A=23-C THUS 4(23-C)+1C=80 OR 92-4C+1C=80 OR -3C=-12 OR C=4 THEN
A+4=23 OR A=23-4 OR A=19 THEN 4*19+1*4=80 OR 76+4=80 OR 80=80